A 5.088 gram sample of an organic compound containing C, H and O is analyzed by combustion analysis and 10.17 grams of CO2 and 4.163 grams of H2O are produced.

In a separate experiment, the molar mass is found to be 88.11 g/mol. Determine the empirical formula and the molecular formula of the organic compound.

  • Enter the elements in the order C, H, O
    empirical formula =
    molecular formula =
Answer #1

any, Let compound is Cx My O2. All. Combution Ron of Cx My Oz. CxHy O2 + 20+ z₂ x Co tyho O mass of Oz = mass of product - ma0.2311 2 X0.2311 0.231) 2 So Empirical formula = Chuo Empirical mass = z 2x12 + 4 + 16 2ut 4+ 16 = 448/mol, n= molar mass Emp